what is long length of right triangle when a=20inches and b=48 inches
then what is the length of c?

Found 2 solutions by solver91311, checkley77:
Answer by solver91311(24713)



Use Pythagoras to find the measure of the hypotenuse of a right triangle given the measures of the two legs.



For your problem:



The arithmetic, as usual, is yours to do. Of course, you could also recognize that 20 is 5 times 4 and 48 is 12 times 4, and recall that 5, 12, and 13 is a Pythagorean Triple hence the hypotenuse of your triangle must be 13 times 4.

John


Answer by checkley77(12844)

A^2+B^2=C^2
20^2+48^2=C^2
400+2,304=C^2
2,704=C^2
C=SQRT2.704
C=52 INCHES ANS. FOR THE LENGTH OF SIDE C.